OpenStudy (anonymous):

how do i write y-5=1/3(x-4) in standard form

OpenStudy (anonymous):

Assuming you meant to write: y -5 = (x- 4)/3 Multiply each side by 3 since fractions are ugly: 3y - 15 = x - 4 Subtract 3y from both sides to move variables to 1 side: -15 = x - 3y - 4 Add 4 to both sides since we want to get rid of 4 from the variable only side. -11 = x - 3y Flip: x - 3y = -11.